if secA-tanA=√3-2. show that (1-sinA)/(1+sinA)=7-4√3

Grade:10

1 Answers

Arun
6 years ago
Dear Shreya
sec A - tan A = sqrt(3) - 2
1- sinA/ cos A = sqrt(3) -2
(1- sinA)/ (1 - sin²A) = sqrt(3) -2
1/1 + sinA = sqrt(3) -2
1 + sinA = 1/[sqrt(3) - 2]
Hope now you can solve it further.
Regards
Arun (askIITians forum expert)